OpenStudy (anonymous):

hello can you help me with some math problems ? The first is solve this differantial equation y´ = -3y, y(0)=1

OpenStudy (anonymous):

given dy/dx =-3y (using method of seperation ) we have dy/y =-3 dx integrating we have ln y -ln c = -3x or ln(y/c) =-3x or y = c e^(-3x) c is integrating constant given at x=0 y=1 hence we have 1= c e^(-3.0) or 1=c hence the required solution is y = e^(-3x)

OpenStudy (anonymous):

Thank you :-)

OpenStudy (anonymous):

welcome....